Grade 8General Physics

2 pistons of hydraulic press have diameters 30cm and 25cm. What is force exerted by large piston, when 50kg weight is placed on smaller piston?

Joshita
7 Years agoGrade 8

1 Answer

Arun
7 Years ago
Dear student
 
ratio of forces =(30/2.5)2=144,
 
force exerted by larger piston=50*144=7200kg wt
 
Regards
Arun